OpenStudy (anonymous):

An object moves in a straight line so that its velocity after t seconds is v(t). Determine the displacement and distance traveled by the object on the given time interval. v(t) = 7 − 14 sin(πt) on [0,2] I found the displacement to be 14. But the distance stinks because of the intervals

OpenStudy (anonymous):

why does the distance stink?

OpenStudy (anonymous):

oh i see, goes backwards then forwards then backwards

OpenStudy (anonymous):

\[14\sin(\pi t)=7\] \[\sin(\pi t)=\frac{1}{2}\] \[\pi t=\frac{\pi}{6}\] \[t=\frac{1}{6}\] for the first one

OpenStudy (anonymous):

ok my first comment was wrong derivative is positive, then negative, then positive so it goes forward then backward, then forward forward on interval \((0,\frac{1}{6})\)

OpenStudy (anonymous):

you good from there? next solve \(\sin(\pi t)=\frac{1}{2}\) \[\pi t=\frac{5\pi}{6}\] \[t=\frac{5}{6}\] and break up your intervals accordingly

OpenStudy (anonymous):

so don't include 1/6?

OpenStudy (anonymous):

the interval is from0->2

OpenStudy (anonymous):

integrate from 0 to 1/6 then from 1/6 tp 5/6 then from 5/6 to 2

OpenStudy (anonymous):

it is negative on \((\frac{1}{6},\frac{5}{6})\) so integrate the negative of the function

OpenStudy (anonymous):

I see. I did that but I guess I need the fraction for rather than the decimal form XP

OpenStudy (anonymous):

you are looking for positive distnace

OpenStudy (anonymous):

right, I got you thanks :)